OpenStudy (anonymous):

Find 7 values of y=5x+2?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i think all you need to do is pick 7 numbers for x and plug them in

OpenStudy (misty1212):

for example if you pick \(x=0\) then you get \(y=0+2=2\) and if you pick \(x=1\) you get \(y=5\times 1+2=7\) and so on

OpenStudy (misty1212):

the truth is of course that \(y\) can be any number, so it is not really clear what it means "find 7 values of y"

OpenStudy (anonymous):

what @misty1212 said y can be anything is that the exact way the question is written or are you supposed to write maybe 7 ordered pairs like \[(0,2),(1,7),(2,12)...\]

OpenStudy (anonymous):

i think seven ordered pairs

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

because it wants me to put it in a table then on a graph

OpenStudy (misty1212):

ok then do what misty said

OpenStudy (misty1212):

lol

OpenStudy (misty1212):

did you get the \((0,2)\) part?

OpenStudy (anonymous):

sorta

OpenStudy (misty1212):

the equation is \[y=5x+2\] right?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

ok so if \(x\) is \(0\) we get \[y=5\times 0+2=2\] right?q

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

that means the ordered pair is \((0,2)\)

OpenStudy (misty1212):

now lets let \(x=1\) we get \(y=5\times 1+2=7\)

OpenStudy (misty1212):

so another ordered pair would be \((1,7)\) ok so far?

OpenStudy (anonymous):

okay

OpenStudy (misty1212):

how about if \(x=2\) what is \(y\) then ?

OpenStudy (anonymous):

12

OpenStudy (misty1212):

ok good now we have 3 ordered pairs \[(0,2),(1,7),(2,12)\] all you need is four more ....

OpenStudy (anonymous):

oh okay i get it wow im slow haha sorry but thank you alot

OpenStudy (misty1212):

lol it is not like you are born knowing it, but you are right, it is not that hard \[\huge \color\magenta\heartsuit\]